weekly wages of 10 workers at random from factory are given below



1
Expert's answer
2022-06-13T14:56:55-0400

We have sample values 578,572,570,568,572,578,570,572,596,584, sample size n=10.

Sample mean

"\bar{x}=\dfrac{1}{10}(578+572+570+568+572"

  


"+578+570+572+596+584)=576"

Sample variance


"s^2=\dfrac{\Sigma(x_i-\bar{x})^2}{n-1}=\dfrac{1}{10-1}(4+16+36+64"




"+16+4+36+16+400+64)=\dfrac{656}{9}""s=\sqrt{s^2}=\sqrt{\dfrac{656}{9}}=\dfrac{4\sqrt{41}}{3}\approx8.5375"


The following null and alternative hypotheses need to be tested:

"H_0:\mu=580"

"H_1:\mu\not=580"

This corresponds to a two-tailed test, for which a t-test for one mean, with unknown population standard deviation, using the sample standard deviation, will be used.

Based on the information provided, the significance level is "\alpha = 0.05," "df=n-1=9" and the critical value for two-tailed test is "t_c =2.262156."

The rejection region for this two-tailed test is "R = \{t:|t|>2.262156\}."

The t-statistic is computed as follows:


"t=\dfrac{\bar{x}-\mu}{s/\sqrt{n}}=\dfrac{576-580}{8.5375/\sqrt{10}}=-1.4816"


Since it is observed that "|t|=1.4816<2.262156=t_c," it is then concluded that the null hypothesis is not rejected.

Using the P-value approach:

The p-value for two-tailed, "df=9" degrees of freedom, "t=-1.4816" is "p=0.172588," and since "p=0.172588>0.05=\alpha," it is concluded that the null hypothesis is not rejected.

Therefore, there is not enough evidence to claim that the population mean "\mu"

is different than 580, at the "\alpha = 0.05" significance level.